# Disordered Dynamics in High Dimensions: Connections to Random Matrices and Machine Learning

### v1

@@ 8.2 Diagonally Modulated Gaussian
This ensemble is of interest in theoretical neuroscience since it describes the eigenvalues of the Jacobian for a randomly connected recurrent neural network sompolinsky1988chaos; helias2020statistical. The DMFT equations for the Hermitianized dynamical system take the form
$i\omega h_{0}(\omega)=-\xi_{2}(\omega)-\sigma^{2}\mathcal{R}_{1}(\omega)h_{0}(\omega)+z^{\star}h_{1}(\omega)$
$h_{1}(\omega)=\xi_{1}+\mathcal{R}_{0,2}(\omega)h_{1}(\omega)-zh_{0}(\omega)+\frac{1}{i\omega}b.$ (151)
We introduce, for each $\sigma$, the response function $\mathcal{H}_{\sigma}(\omega)$ which has the form
@@
From this function, we can compute the averaged response functions $\mathcal{R}_{1}(\omega)$ and $\mathcal{R}_{0,2}(\omega)$ as
@@ 9 Beyond Linear TTI Structure, Evolving Matrices
In the previous examples, the random matrices appearing in the dynamics were frozen. Since the dynamics for $\bm{h}$ were linear, the response functions were time-translation invariant (TTI) and could be obtained directly from a Fourier transform. Subsequently, the correlation functions could be computed in closed form as a two-variable Fourier transform. However, DMFT is arguably most useful in settings beyond linear dynamics. In this section, we illustrate a few simple examples where the resulting DMFT equations remain Gaussian, which enables exact analytical computations of the correlations. A number of interesting high dimensional systems with this Gaussian property have been recently shown to capture similar phenomena such as transitions to chaos, aging dynamics, and potential separations from mean field statics fournier2023statistical; fournier2025non.
@@ 9.1 Toy Example: Simple (Anti)-Hebbian Linear Dynamics
Consider a linear dynamical system with dynamical connectivity $\bm{M}(t)$. While this model was studied in a nonlinear RNN with asymmetric initial connectivity clark2024theory, we will take $\bm{M}(0)$ to be a GOE matrix and study a linear RNN with adaptive weights
@@ 9.2 Deep Linear ResNets under Random Initialization
Another setting where DMFT equations still provide an exact description of the asymptotic dynamics, without having a TTI description is the training dynamics of randomly initialized infinite width neural networks bordelon2022self; bordelon2023depthwise. In this model, all weights of the network start off as random matrices, but experience structured updates from training. In this setting, we will briefly describe a special case of deep linear networks trained on random data in a proportional scaling regime for input dimension $N_{0}$, hidden width $N_{1}$ and dataset size $P$ bordelon_deep_linear
@@
These equations close directly at the level of correlation and response, however the systems dynamics are no longer time-translation invariant so both correlation and response must be solved for simultaneously (unlike the previous examples where response functions could be solved for independently of the correlations). We plot examples of these equations compared to wide but finite networks in Figure 14. Finite $\nu$ and finite $\alpha$ generate corrections to both the bias and variance dynamics of the model. Further if the branch scale parameter $\beta$ is scaled as $\beta=\frac{\beta_{0}}{\sqrt{L}}$, the above dynamics converge to a well defined infinite depth limit as $L\to\infty$. A discrete time version of these equations was explored in bordelon_deep_linear, which enabled comparisons of how optimal learning rate depends on $\nu$ and $L$ in different parameterizations, providing a simple tractable example where the hyperparameter transfer effect can be characterized theoretically yang2021tuning.
@@ 10 Discussion
In this work, we provided a pedagogical overview of DMFT ideas to analyze the evolution of dynamical systems in high dimensions. Our focus was on dynamical systems that admit single site Gaussian processes in the high dimensional limit. However, the potential application areas of DMFT are far more diverse. While we examined the use of DMFT methods for simple models that yield Gaussian processes, many problems generate asymptotic descriptions that are non-Gaussian including descriptions of training dynamics of nonlinear models with random data gerbelot2022rigorous; mignacco2020dynamical; mignacco2022effective, training deep nonlinear networks from random initialization bordelon2022self; bordelon2023depthwise, and Hebbian learning in nonlinear RNNs clark2024theory. In such cases, the single site equations can still be solved with Monte-Carlo sampling to estimate the non-Gaussian single site equations roy2019numerical. Approximations which make the limiting equations as tractable as the examples provided in this work are an active area of research montanari2025dynamical; fournier2023statistical. In general, DMFT is a powerful tool which enables insights into a large variety of complex dynamical systems, including those arising in machine learning theory. We hope that this note inspires future research in this direction.
@@
Figure 14: The dynamics of a deep linear residual network trained with gradient descent. (a) Increasing the parameter $\gamma_{0}$ induces more significant changes in the hidden weights of the network, leading to nonlinear dynamics (the $\gamma_{0}\to 0$ limit results in linear, termed lazy dynamics chizat2019lazy; bordelon2022self). (b) Varying the depth of the residual network converges to a stable predictor if $\beta=\beta_{0}/\sqrt{L}$. (c) Increasing the width to input-dimension ratio $\nu=N_{1}/N_{0}$ reduces the test loss.
@@ Acknowledgements
We thank Clarissa Lauditi for comments on this draft and thank David Clark, Jacob Zavatone-Veth, Alex Atanasov, Courtney Paquette, Elliot Paquette, Francesco Mori, Jingfeng Wu, Yue Lu for insightful discussions. B.B. acknowledges support from the Center of Mathematical Sciences and Applications (CMSA) of Harvard University. C.P. is supported by an NSF CAREER Award (IIS-2239780), DARPA grants DIAL-FP-038 and AIQ-HR00112520041, the Simons Collaboration on the Physics of Learning and Neural Computation, and the William F. Milton Fund from Harvard University. This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence.
@@ Appendix D Structured Random Features
At finite $N,P$ the above path integral is not perfectly dominated by the saddle point. However, the mean field approximation to the test loss is obtained by taking a saddle point over these variables. This incurs an approximation error, but that approximation error decays gracefully as $N,P$ increase [bordelon2020spectrum]. The saddle point equations recover the DMFT equations provided in the main text.
@@ E.1 Asymmetric
As $N\to\infty$, the measure of finitely many vectors $\bm{v}(\omega)$ and $\bm{u}(\omega)$ is uniform up to constraints on the inner products, which is the main insight of low-rank HCIZ integrals [potters2020first]. The inner product constraints on $\bm{v}(\omega)$ and $\bm{u}(\omega)$ are
@@
$\Sigma_{hh}(\omega,\omega^{\prime})=\text{tr}\bm{A}^{2}\left(i\omega+\Psi_{h\chi}(\omega)\bm{A}\right)^{-1}\left[\bm{h}_{0}\bm{h}_{0}^{\top}-\Psi_{\chi\chi}(\omega,\omega^{\prime})\right]\left(i\omega+\Psi_{h\chi}(\omega)\bm{A}\right)^{-1}$
@@ E.2 Symmetrized Case
$\Sigma_{hh}(\omega,\omega^{\prime})=\text{tr}\bm{A}\left(i\omega+\Psi_{h\chi}(\omega)\bm{A}\right)^{-1}\left[\bm{h}_{0}\bm{h}_{0}^{\top}-\bm{A}\Psi_{\chi\chi}(\omega,\omega^{\prime})\right]\left(i\omega+\Psi_{h\chi}(\omega)\bm{A}\right)^{-1}$

### v2

@@ 8.2 Diagonally Modulated Gaussian
This ensemble is of interest in theoretical neuroscience since it describes the eigenvalues of the Jacobian for a randomly connected recurrent neural network Sompolinsky et al. (1988); Helias and Dahmen (2020). The DMFT equations for the Hermitianized dynamical system take the form
$\frac{d}{dt}h_{0}(t)=-h_{2}(t)\ ,\ h_{2}(t)=\xi_{2}(t)+\sigma^{2}\int dt^{\prime}R_{1}(t,t^{\prime})h_{0}(t^{\prime})-z^{\star}h_{1}(t)$
$h_{1}(t)=\xi_{1}(t)+\int dt^{\prime}R_{0,2}(t,t^{\prime})h_{1}(t^{\prime})-zh_{0}(t^{\prime})+b.$ (151)
In the above expressions, the dynamics for $h_{0}$ are modulated by the random variable $\sigma\sim\mu(\sigma)$. The response functions $R_{1}(t,t^{\prime})$ and $R_{0,2}(t,t^{\prime})$ are computed as
$R_{1}(t,t^{\prime})=\left<\frac{\partial h_{1}(t)}{\partial\xi_{1}(t^{\prime})}\right>_{\sigma}\ ,\ R_{0,2}(t,t^{\prime})=\left<\sigma^{2}\frac{\partial h_{0}(t)}{\partial\xi_{2}(t^{\prime})}\right>_{\sigma}$ (152)
where the average $\left$$\cdot\right$> is over the measure $\mu(\sigma)$. To close the equations for the response functions, it will be advantagous to introduce a function $\mathcal{H}_{\sigma}(\omega)\equiv-\int d\tau e^{-i\omega\tau}\frac{\partial h_{0}(t+\tau)}{\partial\xi_{2}(t)}$ which is the response function conditional on a particular value of $\sigma$ .
@@
From this function, we can compute the averaged Fourier-transformed response functions $\mathcal{R}_{1}(\omega)$ and $\mathcal{R}_{0,2}(\omega)$ as
@@ 9 Beyond Linear TTI Structure, Evolving Matrices
In the previous examples, the random matrices appearing in the dynamics were frozen. Since the dynamics for $\bm{h}$ were linear, the response functions were time-translation invariant (TTI) and could be obtained directly from a Fourier transform. Subsequently, the correlation functions could be computed in closed form as a two-variable Fourier transform. However, DMFT is arguably most useful in settings beyond linear dynamics. In this section, we illustrate a few simple examples where the resulting DMFT equations remain Gaussian, which enables exact analytical computations of the correlations. A number of interesting high dimensional systems with this Gaussian property have been recently shown to capture similar phenomena such as transitions to chaos, aging dynamics, and potential separations from mean field statics Fournier and Urbani (2023); Fournier et al. (2025).
@@ 9.1 Toy Example: Simple (Anti)-Hebbian Linear Dynamics
Consider a linear dynamical system with dynamical connectivity $\bm{M}(t)$. While this model was studied in a nonlinear RNN with asymmetric initial connectivity Clark and Abbott (2024), we will take $\bm{M}(0)$ to be a GOE matrix and study a linear RNN with adaptive weights
@@ 9.2 Deep Linear ResNets under Random Initialization
Another setting where DMFT equations still provide an exact description of the asymptotic dynamics, without having a TTI description is the training dynamics of randomly initialized infinite width neural networks Bordelon and Pehlevan (2022); Bordelon et al. (2023). In this model, all weights of the network start off as random matrices, but experience structured updates from training. In this setting, we will briefly describe a special case of deep linear networks trained on random data in a proportional scaling regime for input dimension $N_{0}$, hidden width $N_{1}$ and dataset size $P$ Bordelon and Pehlevan (2025)
@@
These equations close directly at the level of correlation and response, however the systems dynamics are no longer time-translation invariant so both correlation and response must be solved for simultaneously (unlike the previous examples where response functions could be solved for independently of the correlations). We plot examples of these equations compared to wide but finite networks in Figure 14. Finite $\nu$ and finite $\alpha$ generate corrections to both the bias and variance dynamics of the model. Further if the branch scale parameter $\beta$ is scaled as $\beta=\frac{\beta_{0}}{\sqrt{L}}$, the above dynamics converge to a well defined infinite depth limit as $L\to\infty$. A discrete time version of these equations was explored in Bordelon and Pehlevan (2025), which enabled comparisons of how optimal learning rate depends on $\nu$ and $L$ in different parameterizations, providing a simple tractable example where the hyperparameter transfer effect can be characterized theoretically Yang et al. (2021).
@@ 10 Discussion
In this work, we provided a pedagogical overview of DMFT ideas to analyze the evolution of dynamical systems in high dimensions. Our focus was on dynamical systems that admit single site Gaussian processes in the high dimensional limit. However, the potential application areas of DMFT are far more diverse. While we examined the use of DMFT methods for simple models that yield Gaussian processes, many problems generate asymptotic descriptions that are non-Gaussian including descriptions of training dynamics of nonlinear models with random data Gerbelot et al. (2022); Mignacco et al. (2020); Mignacco and Urbani (2022), training deep nonlinear networks from random initialization Bordelon and Pehlevan (2022); Bordelon et al. (2023), and Hebbian learning in nonlinear RNNs Clark and Abbott (2024). In such cases, the single site equations can still be solved with Monte-Carlo sampling to estimate the non-Gaussian single site equations Roy et al. (2019). Approximations which make the limiting equations as tractable as the examples provided in this work are an active area of research Montanari and Urbani (2025); Fournier and Urbani (2023). In general, DMFT is a powerful tool which enables insights into a large variety of complex dynamical systems, including those arising in machine learning theory. We hope that this note inspires future research in this direction.
@@
Figure 14: The dynamics of a deep linear residual network trained with gradient descent. (a) Increasing the parameter $\gamma_{0}$ induces more significant changes in the hidden weights of the network, leading to nonlinear dynamics (the $\gamma_{0}\to 0$ limit results in linear, termed lazy dynamics Chizat et al. (2019); Bordelon and Pehlevan (2022)). (b) Varying the depth of the residual network converges to a stable predictor if $\beta=\beta_{0}/\sqrt{L}$. (c) Increasing the width to input-dimension ratio $\nu=N_{1}/N_{0}$ reduces the test loss.
@@ Acknowledgements
We thank Clarissa Lauditi and Haiping Huang for comments on this draft and thank David Clark, Jacob Zavatone-Veth, Alex Atanasov, Courtney Paquette, Elliot Paquette, Francesco Mori, Jingfeng Wu, Yue Lu for insightful discussions. B.B. acknowledges support from the Center of Mathematical Sciences and Applications (CMSA) of Harvard University. C.P. is supported by an NSF CAREER Award (IIS-2239780), DARPA grants DIAL-FP-038 and AIQ-HR00112520041, the Simons Collaboration on the Physics of Learning and Neural Computation, and the William F. Milton Fund from Harvard University. This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence.
@@ Appendix D Structured Random Features
At finite $N,P$ the above path integral is not perfectly dominated by the saddle point. However, the mean field approximation to the test loss is obtained by taking a saddle point over these variables. This incurs an approximation error, but that approximation error decays gracefully as $N,P$ increase [11]. The saddle point equations recover the DMFT equations provided in the main text.
@@ E.1 Asymmetric
As $N\to\infty$, the measure of finitely many vectors $\bm{v}(\omega)$ and $\bm{u}(\omega)$ is uniform up to constraints on the inner products, which is the main insight of low-rank HCIZ integrals [46]. The inner product constraints on $\bm{v}(\omega)$ and $\bm{u}(\omega)$ are
@@
$\Sigma_{hh}(\omega,\omega^{\prime})=\text{tr}\bm{A}^{2}\left(i\omega+\Psi_{h\chi}(\omega)\bm{A}\right)^{-1}\left[\bm{h}_{0}\bm{h}_{0}^{\top}-\Psi_{\chi\chi}(\omega,\omega^{\prime})\right]\left(i\omega^{\prime}+\Psi_{h\chi}(\omega^{\prime})\bm{A}\right)^{-1}$
@@ E.2 Symmetrized Case
$\Sigma_{hh}(\omega,\omega^{\prime})=\text{tr}\bm{A}\left(i\omega+\Psi_{h\chi}(\omega)\bm{A}\right)^{-1}\left[\bm{h}_{0}\bm{h}_{0}^{\top}-\bm{A}\Psi_{\chi\chi}(\omega,\omega^{\prime})\right]\left(i\omega^{\prime}+\Psi_{h\chi}(\omega^{\prime})\bm{A}\right)^{-1}$